### v1

Abstract
We study the complexity of deterministic and probabilistic inversions of partial computable functions on the reals.
1 Introduction
Turing [26] laid the foundations for a general theory of computation which includes the computability of real functions [1]. Early work [12, 9] cumulated in a robust definition of computable real functions based on effective continuity and a theory of computable analysis [6, 23, 29] which also accommodates the study of computational complexity of real functions [15]. [^1] The RAM-type model of [5] is a different approach and not relevant to the present work. Besides total real functions the theory is just as robust for partial functions, where while loops can restrict their domain [10]. Recent work includes the study of extensions of partial computable real functions to larger domains [14].
The focus of this article is the complexity of inverting total and partial computable functions from reals (infinite binary sequences) to reals.
Despite the vast research in computable analysis, this basic question is largely unexplored, with the exception of elementary facts and the recent work in [13]. This became apparent to us after a question of Levin [17] regarding the existence of partial computable functions that preserve algorithmic randomness and are computationally hard to invert, even probabilistically.
The finite analogues of such oneway functions are fundamental in computational complexity while their existence remains a major open problem [18]. So Levin’s question was whether an answer can be obtained for computable real functions without computational time and space constraints.
An affirmative answer was given by [2] and independently by Gács [11]. However the two results use a different formulations of probabilistic computation.
Toward understanding Levin’s oneway real functions and its variants we study the complexity of inversion in terms of computability focusing on:
  1. (i)
    total and partial computable functions
  2. (ii)
    random-preserving and injective maps
  3. (iii)
    deterministic or (variants of) probabilistic computation
  4. (iv)
    null or positive (measure) domain or range.
According to the standard framework, partial computable real functions $f$ are continuous; the same is true if $f$ is partial computable relative to an oracle. In this fashion we only consider inversions by continuous functions: $g$ inverts $f$ on $y$ if $g(y)$ is in the domain of $f$ and $f(g(y))=y$.
Hardness of inversion could be due to the complexity of the domain or range. Just as the range of a continuous $f$ may not be Borel, the range of a partial computable $f$ may not be arithmetically definable. This aspect is essential and explains the role of measure in the hardness of inversion.
Needed background from computability and randomness along with complexity bounds under (i)-(iv) are given in §2. The role of random oracles (probabilistic inversions) on hardness of inversion is studied in §3 along with the cases where the domain or range is null. We discuss Levin’s notion of oneway functions and compare it with the variant considered by Gács [11].
Under randomness-preservation (in the sense of Martin-Löf [20]) injections are easier to invert. However the existence of an injective partial computable oneway function is unknown. A partial answer is given in §4 where a left-oneway (the notion of Gács [11]) partial computable injection is constructed.
Notation. We use $x,y,\cdots$ to denote reals and $\sigma,\tau,\cdots$ to denote strings. The bit of $x$ at position $n$ is $x(n)$, starting with $n=0$. Let $|\sigma|$ be the length of string $\sigma$. Let $2^{\omega}$ be the collection of all reals and $2^{<\omega}$ the collection of all strings. Let $\preceq$ be the prefix relation between strings and reals, and $\prec$ the strict prefix relation. Let $0^{n}$ be the $n$-bit string of 0s and $0^{\omega}:=000\dots$. For $x,y\in 2^{\omega}$ let
  • $x\upharpoonright_{n}$ be the $n$-bit prefix of $x$ so $x\upharpoonright_{n}=x(0)x(1)\cdots x(n-1)$
  • $x\oplus y$ be the real $z$ with $z(2n)=x(n)$ and $z(2n+1)=y(n)$
A tree $T$ is a non-empty subset of $2^{<\omega}$ such that if $\sigma\prec\tau\in T$ then $\sigma\in T$. A real $x$ is a path of $T$ if $\forall n,\ x\upharpoonright_{n}\in T$ Let $[T]$ be the class of all paths of $T$.
The Cantor space is $2^{\omega}$ with the topology generated by basic open sets
$\llbracket\sigma\rrbracket=\{x\in 2^{\omega}:\sigma\prec x\}$
for $\sigma\in 2^{<\omega}$ which are also clopen. Let $\mu$ be the uniform measure on $2^{\omega}$ defined by $\mu(\llbracket\sigma\rrbracket)=2^{-|\sigma|}$. We identify $2^{\omega}$ with $2^{\omega}\times 2^{\omega}$ via $(x,y)\mapsto x\oplus y$. Classes of reals with positive $\mu$-measure are positive and those with $\mu$-measure 00 are null.
Let $x\leq_{T}z$ denote that $x$ is computable by a Turing machine with oracle $z$. The halting problem relative to $x$ is called the jump of $x$ and is denoted by $x^{\prime}$.
We use the notion of randomness and relative randomness in the sense of Martin-Löf [20] and denote the class of all random reals by $\mathsf{ML}$. We often refer to van Lambalgen’s Theorem [27] (see [8, Corollary 6.9.3]): for everyl $x,y,w$ the real $x\oplus y$ is $w$-random if and only if $x$ is $w$-random and $y$ is $x\oplus w$-random.
2 Complexity bounds and inversions
We give complexity bounds on the domain, range and inversions of partial computable functions, assuming familiarity with the arithmetical, hyperarithmetical and analytic hierarchies on $\mathbb{N},2^{\omega}$. We write $\Pi^{0}_{1}(w)$ for $\Pi^{0}_{1}$ relative to oracle $w$ and similarly for the other definability classes.
2.1 Partial computable functions on the reals
A Turing machine with an infinite oracle tape and an infinite one-way output tape can be regarded as an effective map $f$ on the reals: given $x\in 2^{\omega}$ on its oracle tape the machine runs indefinitely and if the cells of the output tape contain the digits of $y$ we write $y=f(x)$. This is the standard concept of computability on the reals from computable analysis [23, 28]. Let
  • $f:\subseteq 2^{\omega}\to 2^{\omega}$ denote a function from a nonempty subset of $2^{\omega}$ to $2^{\omega}$.
  • $\mathsf{dom}(f)$ be the domain of $f$ and $f(2^{\omega})$ be the range of $f$.
We write $f(x)\downarrow$ if $x\in\mathsf{dom}(f)$ and $f(x)\uparrow$ otherwise. We say that $f$ is partial computable if there is an oracle Turing functional $\Phi$ such that
$\big{(}f(x)\downarrow\iff\forall n,\ \Phi^{x}(n)\downarrow\big{)}\ \land\ \big{(}f(x)\downarrow\implies\forall n,\Phi^{x}(n)=f(x)(n)\big{)}.$
If $\Phi$ has access to an additional oracle $w$ we say that $f$ is partial $w$-computable and write $f\leq_{T}w$. Every $w$-computable function is continuous. Every continuous function with $G_{\delta}$ domain is $w$-computable for some oracle $w$.
A representation is a partial computable $\hat{f}:\subseteq 2^{<\omega}\to 2^{<\omega}$ with
$\big{(}{\sigma\prec\tau\ \wedge\ \hat{f}(\tau)\downarrow}\big{)}\ \implies\ \big{(}{\hat{f}(\sigma)\downarrow\ \wedge\ \hat{f}(\sigma)\preceq\hat{f}(\tau)}\big{)}$
and is a representation of $f:\subseteq 2^{\omega}\to 2^{\omega}$ if $\hat{f}(x\upharpoonright_{n})\prec f(x)$ for all $x,n$ and
$f(x)\downarrow\iff\lim_{n}|\hat{f}(x\upharpoonright_{n})|=\infty.$
A function $f$ is $w$-computable if and only if it has a $w$-computable representation.
2.2 Complexity of domain and range
If $f$ is partial computable its domain is effectively $G_{\delta}$, namely a $\Pi^{0}_{2}$ class. The converse also holds, so the domain of a partial computable $f$ might not be closed.
Proposition 2.1.
If $f:\subseteq 2^{\omega}\to 2^{\omega}$ is partial computable:
  1. (i)
    the domain of $f$ is a $\Pi^{0}_{2}$ class
  2. (ii)
    the restriction of $f$ to any $\Pi^{0}_{2}$ class $P$ is partial computable
  3. (iii)
    every $\Pi^{0}_{2}$ class is the domain of a partial computable $f$.
Proof..
Let $\hat{f}$ be a partial computable representation of $f$. Since
$x\in\mathsf{dom}(f)\iff\forall n\ \exists m\ \exists s\ |\hat{f}_{s}(x\upharpoonright_{m})\downarrow|\geq n$
we get (i). For (ii) let $(V_{i,j})$ be a computable family of clopen sets with
$P=\bigcap_{i}\bigcup_{j}V_{i,j}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptV_{0,0}=2^{\omega},V_{i,j}\subseteq V_{i,j+1}.$
Define the partial computable function $\hat{g}(\sigma)=\hat{f}(\sigma)\upharpoonright_{t_{\sigma}}$ where
$t_{\sigma}:=\max\{\hskip 0.85355pt{i}:{i\leq|\hat{f}(\sigma)|\ \wedge\ \llbracket\sigma\rrbracket\subseteq V_{i,|\sigma|}}\hskip 0.85355pt\}.$
If $\sigma\prec\tau$ and $\hat{g}(\tau)\downarrow$ then $\hat{f}(\tau)\downarrow$ and $\hat{f}(\sigma)\downarrow\preceq\hat{f}(\tau)$ and
$\llbracket\tau\rrbracket\subset\llbracket\sigma\rrbracket\ \implies\ t_{\tau}\geq t_{\sigma}\ \implies\ \hat{g}(\sigma)\downarrow\preceq\hat{g}(\tau)$
so $\hat{g}$ is a representation of a partial computable $g$. If $x\in\mathsf{dom}(f)\cap P$ then
$\forall i\ \exists k,j,\ \llbracket x\upharpoonright_{k}\rrbracket\subseteq V_{i,j}.$
If $n\geq k,j$, $|\hat{f}(x\upharpoonright_{n})|\geq i$ by $\llbracket x\upharpoonright_{n}\rrbracket\subseteq V_{i,n}$ we get $|\hat{g}(x\upharpoonright_{n})|\geq i$. Such $n$ exists for all $i$ so $g(x)\downarrow$. Since $\forall n,\ \hat{g}(x\upharpoonright_{n})\preceq\hat{f}(x\upharpoonright_{n})\prec f(x)$ we get $g(x)=f(x)$.
If $x\notin\mathsf{dom}(f)$ then $g(x)\uparrow$. If $x\notin P$ then
$\exists i\ \forall j,\ x\notin V_{i,j}\ \implies\ \forall n,\ |\hat{g}(x\upharpoonright_{n})|\leq i\ \implies\ g(x)\uparrow.$
So $g$ is the restriction of $f$ to $\mathsf{dom}(f)\cap P$. Finally (iii) follows from (ii). ∎
Continuous functions $f$ on subsets of $2^{\omega}$ map compact sets to compact sets. Since $2^{\omega}$ is a Hausdorff space, the same holds for closed sets. An effective version of this fact is true when $f$ is partial computable.
Recall that if $P\in\Pi^{0}_{1}(w)$ there is a $w$-computable tree $\hat{P}$ with $P=[\hat{P}]$.
Proposition 2.2.
Let $f:\subseteq 2^{\omega}\to 2^{\omega}$ be partial computable. For any $w\in 2^{\omega}$:
  1. (i)
    if $P\subseteq\mathsf{dom}(f)$ then $P\in\Pi^{0}_{1}(w)\ \implies\ f(P)\in\Pi^{0}_{1}(w)$,
  2. (ii)
    if $Q\in\Pi^{0}_{1}(w)$ there is $P\in\Pi^{0}_{1}(w)$ with $f^{-1}(Q)=\mathsf{dom}(f)\cap P$.
Proof..
For (i) let $\hat{P}$ be a $w$-computable tree with $[\hat{P}]=P$. We show
$y\in f(P)\iff\forall n\ \exists\sigma\in\hat{P}\cap 2^{n},\ \hat{f}(\sigma)\prec y.$ (1)
If $y\in f(P)$ then $\exists x\in P,\ f(x)=y$ so $\forall n\ \big{(}{x\upharpoonright_{n}\in\hat{P}\cap 2^{n}\ \wedge\ \hat{f}(x\upharpoonright_{n})\prec y}\big{)}$.
For the converse, suppose that there is $\sigma_{0},\sigma_{1},\dots$ with $\sigma_{i}\in\hat{P}\cap 2^{i}$ and $\hat{f}(\sigma_{i})\prec y$. Since $2^{\omega}$ is compact, there is a subsequence of $\sigma_{i}$ (without loss of generality, the sequence $\sigma_{i}$ itself) that converges to some $x$. Let $\tau_{i}$ be the longest common prefix of $\sigma_{i}$, $x$. Since $\sigma_{i}\to x$, we get $|\tau_{i}|\to\infty$ so $\tau_{i}\to x$. Also
  • since $\sigma_{i}\in\hat{P}$ we get $\tau_{i}\in\hat{P}$ so $x\in[\hat{P}]=P\subseteq\mathsf{dom}(f)$ and $f(x)\downarrow$
  • since $\hat{f}(\tau_{i})\preceq\hat{f}(\sigma_{i})\prec y$ and $\tau_{i}\to x$ we get $f(x)=y$.
This completes the proof of (1) so $f(P)\in\Pi^{0}_{1}(w)$. For (ii) we have
$x\in f^{-1}(Q)\iff x\in\mathsf{dom}(f)\ \land\ \forall n\ \forall i\ \ \llbracket\hat{f}(x\upharpoonright_{n})\rrbracket\cap Q_{i}\neq\emptyset$
where each $Q_{i}$ is clopen, $i\mapsto Q_{i}$ is $w$-computable and $Q=\bigcap_{i}Q_{i}$. Then
$P=\{x:\forall n\ \forall i\ \llbracket\hat{f}(x\upharpoonright_{n})\rrbracket\cap Q_{i}\neq\emptyset\}$
is $\Pi^{0}_{1}(w)$ and $f^{-1}(Q)=\mathsf{dom}(f)\cap P$. ∎
Although the domain of any partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ is arithmetically definable (indeed $\Pi^{0}_{2}$) as a subset of $2^{\omega}$ it is possible that it is nonempty and has no arithmetically definable member.
Proposition 2.3.
The following hold for $f:\subseteq 2^{\omega}\to 2^{\omega}$:
  1. (i)
    if $x\in\Delta^{1}_{1}$ there is a partial computable $f$ with domain $\{z\}$ and $x\leq_{T}z$
  2. (ii)
    the domain of every partial computable $f$ contains a real in $\Delta^{1}_{1}$.
Proof..
For each $x\in\Delta^{1}_{1}$ there is an ordinal $\alpha<\omega_{1}^{\rm CK}$ such that $x\leq_{T}\emptyset^{(\alpha)}$. By [7, Theorem 2.1.4] $\{\emptyset^{(\alpha)}\}$ is $\Pi^{0}_{2}$. So (i) holds by Proposition 2.1 (iii). By the Gandy Basis Theorem [24, III.1.5] every nonempty $\Sigma^{1}_{1}$ class has a $\Delta^{1}_{1}$ member. So (ii) holds by Proposition 2.1 (i). ∎
In particular a partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ may have a non-empty domain such that all arithmetically definable closed subsets (represented as the set of infinite paths through an arithmetically definable tree) of it are empty.
By Proposition 2.2 (i) the range of any total computable $f:2^{\omega}\to 2^{\omega}$ is $\Pi^{0}_{1}$. On the other hand the range of a partial computable $f$ is only $\Sigma^{1}_{1}$. The latter follows by the representation of $\Sigma^{1}_{1}$ classes with quantifiers over the Baire space $\omega^{\omega}$ and an effective embedding of $\omega^{\omega}$ into $2^{\omega}$.
Proposition 2.4.
The following hold:
  1. (i)
    the range of any partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ is $\Sigma^{1}_{1}$
  2. (ii)
    each $\Sigma^{1}_{1}$ class is the range of a partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$.
Proof..
Clause (i) is straightforward. For (ii) let $Q\in\Sigma^{1}_{1}$.
There exists a total Turing functional $\Phi$ such that
  • $\forall\alpha\in\omega^{\omega}$ there exists at most one $y\in 2^{\omega}$ with $\forall s,\ \Phi(\alpha\oplus y)(s)=1$
  • $y\in Q\iff\exists\alpha\in\omega^{\omega}\ \ \forall s:\ \Phi(\alpha\oplus y)(s)=1$.
Since $\Phi$ is total the function $\varphi:\subseteq\omega^{\omega}\to 2^{\omega}$ given by
$\varphi(\alpha)=y\iff\forall s:\ \Phi(\alpha\oplus y)(s)=1$
is partial computable. Define an effective injection from $\omega^{\omega}$ to $2^{\omega}$ by
$\pi(\alpha):=0^{1+\alpha(0)}10^{1+\alpha(1)}\dots 0^{1+\alpha(n)}\dots$
Then $f(x):=\varphi(\pi^{-1}(x))$ is partial computable and $f(2^{\omega})=Q$. ∎
When $f$ has positive domain it has an effective restriction with $\Pi^{0}_{1}(\emptyset^{\prime})$ range.
Proposition 2.5.
If $f:\subseteq 2^{\omega}\to 2^{\omega}$ is partial computable with positive domain
  1. (i)
    there is a positive $P\in\Pi^{0}_{1}(\emptyset^{\prime})$, $P\subseteq\mathsf{dom}(f)$ with $f(P)\in\Pi^{0}_{1}(\emptyset^{\prime})$
  2. (ii)
    the restriction of $f$ to $P$ is partial computable
and $P$ consists entirely of 2-random reals.
Proof..
By Kurtz [16] (see [8, Theorem 6.8.3]) every positive $\Pi^{0}_{2}$ class has a positive $\Pi^{0}_{1}(\emptyset^{\prime})$ subclass $P$. By intersecting $P$ with a sufficiently large $\Pi^{0}_{1}(\emptyset^{\prime})$ class of 2-randoms $P$ can be chosen to be a class of 2-randoms. So (i) follows by Proposition 2.1 (i) and Proposition 2.2 (i). Proposition 2.1 (ii) gives (ii). ∎
By Proposition 2.5 the complexity bound in Proposition 2.4 can be reduced by considering partial computable restrictions of $f$.
2.3 Complexity of inversion
If $f,g:\subseteq 2^{\omega}\to 2^{\omega}$ we say $g$ is an inversion of $f$ if $\forall y\in f(2^{\omega}),\ f(g(y))=y$.
By [2, Corollary 3.2] every total computable injection $f$ has a total computable inversion. We are particularly interested in functions that preserve randomness.
Definition 2.6.
We say that $f:\subseteq 2^{\omega}\to 2^{\omega}$ is random-preserving if it has positive domain and $f(x)$ is random for each random $x\in\mathsf{dom}(f)$.
By [21, Theorem 4.3], if $x$ is $w$-random and $y\leq_{T}x$ is random then $y$ is $w$-random. So partial computable random-preserving functions preserve $w$-randomness for all $w$.
Theorem 2.7.
There is a total computable random-preserving surjection $f$ which has no continuous inversion.
Proof..
Let $(z^{\prime}_{s})$ be the universal enumeration of the jump $z^{\prime}$ of $z$ and set
$p^{z}(\langle n,s\rangle)$ :={ $:=\ \begin{cases}\ \ 2n&\textrm{if $n\in z^{\prime}_{s+1}-z^{\prime}_{s}$}\\[5.69046pt] 2\langle n,s\rangle+1&\text{otherwise}\end{cases}$
$h^{z}(x)(\langle n,s\rangle)$ $:=\ x(p^{z}(\langle n,s\rangle))$
By the relativization of [2, Theorem 4.5]:
  1. (i)
    $h^{z}$ is a $z$-computable $z$-random preserving total surjection
  2. (ii)
    if $w$ computes an inversion of $h^{z}$ then $w\geq_{T}z^{\prime}$.
Since $z\mapsto h^{z}$ is effective, by [2, Lemma 5.5] the map
$f(x\oplus z):=h^{z}(x)\oplus z$
is a total computable random-preserving surjection. For contradiction, assume that $g$ is a continuous inversion of $f$. Without loss of generality we suppose that $\mathsf{dom}(g)=f(2^{\omega})$ which is $\Pi^{0}_{2}$, so $g$ has $G_{\delta}$ domain and therefore must be $w$-computable for some $w$. Now
$\forall y,z:\ f(g(y\oplus z))=y\oplus z.$
Let $g^{z}(y)$ be the even bits of $g(y\oplus z)$ so $g^{z}$ is an inversion of $h^{z}$ and
$g^{z}\leq_{T}z\oplus w\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\forall z,\ z^{\prime}\leq_{T}z\oplus w$
by (ii). In particular $w^{\prime}\leq_{T}w\oplus w$ which is a contradiction. ∎
Partial inversions of effective maps on the reals are more tractable.
Definition 2.8.
Given $f,g:\subseteq 2^{\omega}\to 2^{\omega}$ we say that $g$ is a
  • positive inversion of $f$ if $\mu(\{\hskip 0.85355pt{y}:{f(g(y))=y}\hskip 0.85355pt\})>0$.
  • probabilistic inversion of $f$ if $\mu(\{\hskip 0.85355pt{y\oplus r}:{f(g(y\oplus r))=y}\hskip 0.85355pt\})>0$.
By [2, Theorem 4.5] there is a total computable random-preserving surjection $f$ such that every oracle that computes a probabilistic inversion of $f$ computes $\textup{{0}}^{\prime}$. The converse is:
Theorem 2.9.
Each total computable $f:2^{\omega}\to 2^{\omega}$ with positive range has a $\textup{{0}}^{\prime}$-computable positive inversion.
Proof..
Clearly $f^{-1}(y)\in\Pi^{0}_{1}(y)$ uniformly in $y$: $\{\hskip 0.85355pt{x\oplus y}:{x\in f^{-1}(y)}\hskip 0.85355pt\}\in\Pi^{0}_{1}$.
So there exists a Turing functional $\Phi$ such that
$\forall y\in f(2^{\omega}):\ \Phi(y^{\prime})\in f^{-1}(y).$
By [8, Theorem 8.14.5] we have $y^{\prime}\leq_{T}y\oplus\emptyset^{\prime}$ for almost every $y$. Since $f(2^{\omega})$ is positive, by countable additivity there exists a Turing functional $\Psi$ with
[…]
Let $g(y)=\Phi(\Psi(y\oplus\emptyset^{\prime}))$ so $g\leq_{T}\textup{{0}}^{\prime}$ and $g$ inverts $f$ on
$\{\hskip 0.85355pt{y\in f(2^{\omega})}:{g(y)\in f^{-1}(y)}\hskip 0.85355pt\}=\{\hskip 0.85355pt{y}:{f(g(y))=y}\hskip 0.85355pt\}\supseteq L.$
Since $\mu(L)>0$ this is a positive subset of $f(2^{\omega})$. ∎
We can now characterize the complexity of positive and probabilistic inversions of total computable functions $f:2^{\omega}\to 2^{\omega}$ with positive range.
Corollary 2.10.
The following are equivalent for $w\in 2^{\omega}$:
  1. (i)
    $w$ computes a positive inversion of every total computable $f$
  2. (ii)
    $w$ computes a probabilistic inversion of every total computable $f$
  3. (iii)
    $w$ computes $\textup{{0}}^{\prime}$
where $f:2^{\omega}\to 2^{\omega}$ is assumed to have positive range.
Proof..
By [2, Theorem 4.5] mentioned above we get (ii)$\to$(iii) while (i)$\to$(ii) is immediate. Theorem 2.9 gives (iii)$\to$(i). ∎
It is not hard to show that $\textup{{0}}^{\prime}$ does not compute a positive inversion of some partial computable $f:2^{\omega}\to 2^{\omega}$ with positive range. We show that under a mild condition on $f$ oracle $\textup{{0}}^{\prime\prime}$ is sufficient for this purpose. The condition is that $f$ maps to random reals with positive probability, which is weaker than randomness-preservation and present in Levin’s oneway functions [2].
Theorem 2.11.
Every partial computable $f$ with $\mu(f^{-1}(\mathsf{ML}))>0$ has a positive inversion which is computable in $\textup{{0}}^{\prime\prime}$.
Proof..
By Proposition 2.5 (ii) we may assume that $\mathsf{dom}(f)\in\Pi^{0}_{1}(\emptyset^{\prime})$. Then
$\{\hskip 0.85355pt{x\oplus y}:{x\in f^{-1}(y)}\hskip 0.85355pt\}\in\Pi^{0}_{1}(\emptyset^{\prime})$ (2)
so $f^{-1}(y)\in\Pi^{0}_{1}(y\oplus\emptyset^{\prime})$ uniformly in $y$ and for some Turing functional $\Phi$:
$\forall y\in f(2^{\omega}):\ \Phi((y\oplus\emptyset^{\prime})^{\prime})\in f^{-1}(y).$
By [8, Theorem 8.14.5] for almost all $y$ we have $(y\oplus\emptyset^{\prime})^{\prime}\equiv_{T}y^{\prime\prime}\equiv_{T}y\oplus\emptyset^{\prime\prime}$.
Since $\mu(f^{-1}(\mathsf{ML}))>0$ there exists a 2-random $x$ such that $f(x)$ is random. By [21, Theorem 4.3] and $f(x)\leq_{T}x$ we get that $f(x)$ is a 2-random.
Since $f(2^{\omega})\in\Pi^{0}_{1}(\emptyset^{\prime})$ we have $\mu(f(2^{\omega}))>0$ and by countable additivity there exists a Turing functional $\Psi$ with $\mu(L)>0$ where
$L:=\{\hskip 0.85355pt{y\in f(2^{\omega})}:{\Psi(y\oplus\emptyset^{\prime\prime})=(y\oplus\emptyset^{\prime})^{\prime}}\hskip 0.85355pt\}.$
Let $g(y)=\Phi(\Psi(y\oplus\emptyset^{\prime\prime}))$ so $g\leq_{T}\textup{{0}}^{\prime\prime}$ and $g$ inverts $f$ on
$\{\hskip 0.85355pt{y\in f(2^{\omega})}:{g(y)\in f^{-1}(y)}\hskip 0.85355pt\}=\{\hskip 0.85355pt{y}:{f(g(y))=y}\hskip 0.85355pt\}\supseteq L.$
Since $\mu(L)>0$ this is a positive subset of $f(2^{\omega})$. ∎
3 Deterministic and probabilistic inversion
A partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ is nowhere effectively invertible if
$x\not\leq_{T}f(x)$   for all $x\in\mathsf{dom}(f)$. (3)
In §3.1 we show that this basic hardness for inversion condition is not hard to satisfy, even in the presence of additional requirements. However these functions are invertible with the help of a random oracle, in certain ways. Hardness with respect to probabilistic inversions is explored in §3.2.
3.1 Inversion without random oracles
The simplest example of a nowhere effectively invertible partial computable function is the projection $f:\ x\oplus z\mapsto x$ restricted to a positive $\Pi^{0}_{1}$ class $P$ of randoms. However $f$ is a.e.  probabilistically invertible in the sense that
$\mu(\{\hskip 0.85355pt{y\oplus r}:{f(g(y\oplus r))=y}\hskip 0.85355pt\})=\mu(f(2^{\omega}))>0$ (4)
for some partial computable $g$.
Proposition 3.1.
There is a partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ which is
  1. (i)
    random-preserving and nowhere effectively invertible
  2. (ii)
    a.e.  probabilistically invertible.
Proof..
Let $f$ be the restriction of $x\oplus z\mapsto x$ to a positive $\Pi^{0}_{1}$ class $P$ of randoms. Then $f$ is partial computable, random-preserving and
$x\oplus z\in P\ \implies\ x\oplus z\not\leq_{T}x$
by van Lambalgen’s theorem, so $f$ is nowhere effectively invertible. By Fubini’s theorem (4) holds for $g(y\oplus r):=y\oplus r$ so (ii) holds. ∎
We turn to examples with null (and nonempty) domain or range.
Proposition 3.2.
There is a partial computable injection $f:\subseteq 2^{\omega}\to 2^{\omega}$ such that no arithmetically definable $g$ can invert $f$ on any real.
Proof..
By Proposition 2.3 (i) $\{\emptyset^{(\omega)}\}\in\Pi^{0}_{2}$ so by Proposition 2.1 (ii)
$f:\emptyset^{(\omega)}\mapsto 0^{\omega}\hskip 8.5359pt\textrm{is partial computable.}$.
Any $g$ that inverts $f$ on $y=0^{\omega}$ defines $\emptyset^{(\omega)}$ which is not arithmetical. ∎
A nowhere effectively invertible partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ with
positive  $\mathsf{dom}(f)\subseteq\mathsf{ML}$    and    $f(2^{\omega})\cap\mathsf{ML}=\emptyset$
can be given by Kurtz [16] (see [8, Theorem 8.21.3]) who constructed $f$ with the above properties and every real in its range is 1-generic. Since 1-generics do not compute randoms [8, Theorem 8.20.5], $f$ is nowhere effectively invertible.
We end with the symmetric case of null domain and positive range.
Proposition 3.3.
There is a partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ which is
  1. (i)
    nowhere effectively invertible
  2. (ii)
    $\mathsf{dom}(f)\cap\mathsf{ML}=\emptyset$ and $f(2^{\omega})\subseteq\mathsf{ML}$ and $\mu(f(2^{\omega}))>0$.
Proof..
Let $Q\neq\emptyset$ be a positive $\Pi^{0}_{2}$ class of 2-randoms and
$f(x\oplus z):\simeq\begin{cases}z&\textrm{if $x=\emptyset^{\prime}$ and $z\in Q$}\\ \uparrow&\textrm{otherwise.}\end{cases}$
Since $x=\emptyset^{\prime}$ is $\Pi^{0}_{2}$ by Proposition 2.1 (ii) $f$ is partial computable. By the choice of $Q$ we get (ii). Since 2-randoms do not compute $\emptyset^{\prime}$ we get (i). ∎
3.2 Inversion with random oracles
In computational complexity oneway functions are effectively calculable finite maps that cannot be effectively inverted with positive probability. Levin [17] defined oneway functions $f:\subseteq 2^{\omega}\to 2^{\omega}$ to be partial computable with
$\mu(f^{-1}(\mathsf{ML}))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu(\{\hskip 0.85355pt{y\oplus r}:{f(g(y\oplus r))=y}\hskip 0.85355pt\})=0$ (5)
for each partial computable $g$. A total computable oneway function was constructed in [2]. Gács [11] constructed a partial computable $f$ with
$\mu(\mathsf{dom}(f))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu\big{(}{\{\hskip 0.85355pt{x\oplus r}:{f(g(f(x),r))=f(x)}\hskip 0.85355pt\}}\big{)}=0$
for each partial computable $g$. This differs from (5) in that probability refers to the domain rather than the range of $f$. Another difference is condition $\mu(f^{-1}(\mathsf{ML}))>0$ which is weaker than randomness preservation but stronger than $\mu(\mathsf{dom}(f))>0$. This suggests two variants of probabilistic inversion.
Definition 3.4.
Let $f:\subseteq 2^{\omega}\to 2^{\omega}$ be partial computable. If
$\mu(\mathsf{dom}(f))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu\big{(}{\{\hskip 0.85355pt{x\oplus r}:{f(g(f(x),r))=f(x)}\hskip 0.85355pt\}}\big{)}=0.$
for each partial computable $g$ we say that $f$ is left-oneway. If
$\mu(f(2^{\omega}))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu(\{y\oplus r:f(g(y\oplus r))=y\})=0$
for each partial computable $g$ we say that $f$ is right-oneway.
Given the absence of $\mu(f^{-1}(\mathsf{ML}))>0$ in the above definition note that
$\mu(f^{-1}(\mathsf{ML}))>0\underset{\not\Leftarrow}{\Rightarrow}\big{(}{\mu(\mathsf{dom}(f))>0\ \wedge\ \mu(f(2^{\omega}))>0}\big{)}.$ (6)
The first implication is trivial and by Proposition 2.5 we may assume that the domain and range of $f$ are $\Pi^{0}_{1}(\emptyset^{\prime})$. Since $\mu(f^{-1}(\mathsf{ML}))>0$ there is a 2-random $x$ such that $f(x)$ is random. By [21, Theorem 4.3] we get that $f(x)$ is a 2-random member of $f(2^{\omega})$ which is $\Pi^{0}_{1}(\emptyset^{\prime})$ so $\mu(f(2^{\omega}))>0$.
The failure of the converse in (6) is demonstrated later in Theorem 3.8.
Before we separate the classes of left and right oneway functions we show that $\mu(f^{-1}(\mathsf{ML}))>0$ is equivalent (up to restrictions) to randomness-preservation.
Lemma 3.5.
If $f:\subseteq 2^{\omega}\to 2^{\omega}$ is partial computable then $\mu(f^{-1}(\mathsf{ML}))>0$ if and only if $f$ extends a random-preserving partial computable function.
Proof..
Suppose that $f$ extends a random-preserving partial computable $h$. Then $h$ has a positive domain $P$ and by Proposition 2.5 we may assume that $P$ consists of randoms. Since $h$ is random-preserving $h^{-1}(\mathsf{ML})=P$ so
$\mu(f^{-1}(\mathsf{ML}))\geq\mu(h^{-1}(\mathsf{ML}))=\mu(P)>0.$
For the converse assume that $\mu(f^{-1}(\mathsf{ML}))>0$ so there is
  • $Q\in\Pi^{0}_{1}$ of randoms with $\mu(f^{-1}(Q))>0$
  • $P\in\Pi^{0}_{1}$ of randoms with $\mu(P\cap f^{-1}(Q))>0$.
Let $g$ be the restriction $f$ to $P\cap f^{-1}(Q)$. The by Proposition 2.2 (ii) the class $P\cap f^{-1}(Q)$ is $\Pi^{0}_{2}$ so $g$ is partial computable by Proposition 2.1 (ii). Since $P\cap f^{-1}(Q)$ is positive and $P,Q$ consist of randoms, $g$ is random-preserving. ∎
Assuming randomness-preservation right-oneway implies left-oneway.
Theorem 3.6.
Every partial computable random-preserving right-oneway function $f$ is left-oneway.
Proof..
Suppose that $f$ is random-preserving so it has positive domain.
If $f$ is not left-oneway there is a partial computable function $g$ with
$\mu(L)>0\hskip 8.5359pt\textrm{where}\hskip 8.5359ptL:=\{x\oplus r:f(g(f(x),r))=f(x)\}$
To show that $f$ is not right-oneway it suffices to show that
$\mu(R)>0\hskip 8.5359pt\textrm{where}\hskip 8.5359ptR:=\{y\oplus r:f(g(y,r))=y\}.$
Since $\mu(L)>0$ there is a $2$-random $x\oplus r\in L$. By van Lambalgen’s theorem $x$ is $r\oplus\emptyset^{\prime}$ random. Since $f$ is random-preserving $f(x)$ is random and since $f(x)\leq_{T}x$ by [21, Theorem 4.3] it is also $r\oplus\emptyset^{\prime}$-random. By van Lambalgen’s theorem $f(x)\oplus r$ is a $2$-random member of the $\Pi^{0}_{2}$ class $R$ so $\mu(R)>0$. ∎
The converse is not true, even for total computable random-preserving $f$.
Theorem 3.7.
There is a random-preserving partial computable left-oneway $f$ which is not right-oneway.
Proof..
Let $f$ be a random-preserving partial computable oneway function, which exists by [2, Theorem 4.4]. Let $\bar{f}$ be defined by
$\bar{f}(x)=\begin{cases}0f(z)&\text{ if }x=0z\text{ for some }z\\ 1z&\text{ if }x=1(0^{\omega}\oplus z)\text{ for some }z\\ \ \uparrow&\text{ otherwise}\end{cases}$
then $\bar{f}$ is random-preserving since $f$ is. The partial computable function
$1z\mapsto 1(0^{\omega}\oplus z)$
inverts $\bar{f}$ on all $y\in\llbracket 1\rrbracket$, so $\bar{f}$ is not one-way or right-oneway.
One the other hand, any partial computable $\bar{g}$ such that
$\mu(\{x:\bar{f}(\bar{g}(\bar{f}(x),r)=\bar{f}(x)\})>0$
induces a partial computable $g$ such that
$\mu(\{x:f(g(f(x),r))=f(x)\})>0$
which is impossible since $f$ is left-oneway. Therefore $\bar{f}$ is left-oneway. ∎
We show that Theorem 3.6 fails if we do not require randomness-preservation. We say that a real $x$ is incomplete if $\emptyset^{\prime}\not\leq_{T}x$.
Theorem 3.8.
There is a total computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ such that:
  • $f$ is a surjection with $f(\mathsf{ML})\cap\mathsf{ML}=\emptyset$
  • $f$ is right-oneway but not left-oneway.
Proof..
Let $P=\bigcap_{s}P_{s}$ be the $\Pi^{0}_{1}$ class of $\mathsf{PA}$ reals where $P_{s}$ is clopen and $s\mapsto P_{s}$ is computable. Define the total computable surjection $f:2^{\omega}\to 2^{\omega}$ by
$f(x\oplus z)=\begin{cases}\ \ z&\textrm{if $x\in P$}\\ z\upharpoonright_{n}\ 0^{\omega}&\textrm{if $n=\min\{\hskip 0.85355pt{s}:{x\not\in P_{s}}\hskip 0.85355pt\}$.}\end{cases}$
If $x\oplus z$ is random, $x\not\in P$ so $f(x\oplus z)$ is computable. So $f(\mathsf{ML})\cap\mathsf{ML}=\emptyset$.
Almost all reals $y\in f(2^{\omega})$ are incomplete randoms. For each random $y$ every $w\in f^{-1}(y)$ computes a $\mathsf{PA}$ real. So for almost all $r$ and $y\in f(2^{\omega})$:
  • $y\oplus r$ is an incomplete random
  • each real in $f^{-1}(y)$ computes a $\mathsf{PA}$ real.
By [25] incomplete randoms do not compute $\mathsf{PA}$ reals so
$\mu(\{\hskip 0.85355pt{r}:{\textrm{$r\oplus y$ computes a member of $f^{-1}(y)$}}\hskip 0.85355pt\})=0$
for almost all $y\in f(2^{\omega})$. This shows that $f$ is right-oneway. Since
$x\not\in P\implies f(x\oplus z)\ \text{ends with}\ 0^{\omega}$
the set $f((2^{\omega}-P)\times 2^{\omega})$ is countable. Since $\mu(P)=0$ by countable additivity there is $y$ with $\mu(f^{-1}(y))>0$. For this $y$ we have
$\mu(\{v\oplus r:f(v)=f(r)=y\})=\mu(f^{-1}(y))^{2}>0$
so for the function $g(w,r):=r$ we get
$\mu\left(\{v\oplus r:f(g(f(v),r))=f(v)\}\right)>0.$
Since $g$ is computable this shows that $f$ is not left-oneway. ∎
4 A left-oneway injection
Virtually all examples of partial computable functions that are hard to invert we have seen are non-injective, with the exception of Proposition 3.2. Random oracles do not reduce the complexity of inverting injective maps.
Lemma 4.1.
Let $f:\subseteq 2^{\omega}\to 2^{\omega}$ be a partial computable injection with positive domain. If $x\not\leq_{T}f(x)$ for almost all $x\in\mathsf{dom}(f)$ then $f$ is left-oneway.
Proof..
Since $f$ is injective we have
$\{x\oplus r:f(g(f(x)\oplus r))=f(x)\}=\{x\oplus r:g(f(x)\oplus r)=x\}.$
For a contradiction assume that $f$ is not left-oneway so
$\mu(\{x\oplus r:g(f(x)\oplus r)=x\})>0$
for some partial computable $g$. Then there is $x\in\mathsf{dom}(f)$ with $x\not\leq_{T}f(x)$ and
$\mu(\{r:g(f(x)\oplus r)=x\})>0$
so $\mu(\{r:x\leq_{T}f(x)\oplus r\})>0$ which contradicts $x\not\leq_{T}f(x)$. ∎
This raises the question whether all effective injections are easy to invert.
Theorem 4.2.
There is a left-oneway partial computable injection.
The proof is an effective measure-theoretic argument in a well-established framework for the construction of effective probabilistic maps [3]. It is of independent interest since similar results [19, 4, 22] do not give injective maps.
4.1 Outline of the proof
Our argument is self-contained but follows the standard framework for constructing probabilistic maps, see [8, §8.21] or [3, 4].
Let $(\Phi_{e})$ be an effective enumeration of all Turing functionals. We construct a partial computable injectiion $f$ with positive domain such that for all $e$:
$R_{e}\text{:}\ \ \mu(\{x:\Phi_{e}(f(x))\downarrow=x\})=0.$
We construct a representation $\hat{f}$ of $f$ in stages. At stage $s$ the domain of the current approximation $\hat{f}_{s}$ of $\hat{f}$ is a finite tree where each node with a successor has two successors. The leaf nodes of $\hat{f}_{s}$ are the leaf nodes of its domain.
Satisfying $R_{e}$. Suppose that $\hat{f}(\sigma)$ is defined and all $\hat{f}(\sigma\tau),\tau\neq\lambda$ are undefined. To satisfy $R_{e}$ in $\llbracket\sigma\rrbracket$ we need to define $f(x)$ so that for almost all $x\in\llbracket\sigma\rrbracket$
$\Phi_{e}(f(x))\uparrow\ \ \vee\ \ \ \Phi_{e}(f(x))\neq x.$
Case 1. If $\Phi_{e}(y)\uparrow$ for all $y\in f(\llbracket\sigma\rrbracket)$ then $R_{e}$ is automatically satisfied.
Case 2. If $\Phi_{e}(y)(|\sigma|)\downarrow=i$ for some $y\in f(\llbracket\sigma\rrbracket)$ with oracle-use $\rho\prec y$ then
$\big{(}{\rho\prec f(x)\ \wedge\ x(|\sigma|)\neq i}\big{)}\ \implies\ \Phi_{e}(f(x))(|\sigma|)=\Phi_{e}(y)(|\sigma|)\neq x(|\sigma|)$ (7)
which means that $R_{e}$ is satisfied for this $x$. So we
  1. (a)
    satisfy $R_{e}$ in $\llbracket\sigma(1-i)\rrbracket$ by defining $\hat{f}(\sigma(1-i))=\rho$
  2. (b)
    define $\hat{f}(\sigma i)$ arbitrarily and try to satisfy $R_{e}$ with $\sigma i$ in place of $\sigma$.
By iterating we satisfy $R_{e}$ for almost all $x\in\llbracket\sigma\rrbracket$.
Non-uniformity. Unfortunately we cannot effectively distinguish which case is true. We deal with this by creating a branch $\sigma\tau$ above $\sigma$ for each $\tau$ of certain length $l+1$ and assign to them different guesses. Branches $\sigma 00^{l}$, $\sigma 10^{l}$:
  • believe case 2 and search for $\rho$ as in (7)
  • if $\rho$ is found they define $\hat{f}(\sigma(1-i)0^{l})\succ\rho$ satisfying $R_{e}$ in $\llbracket\sigma(1-i)0^{l}\rrbracket$.
The $\hat{f}$-value at $\sigma i0^{l}$ is defined arbitrarily and $R_{e}$ in $\llbracket\sigma i0^{l}\rrbracket$ is considered by the nodes above. If the guess is wrong:
  • the search never stops and $f(x)\uparrow$ for $x\in\llbracket\sigma 00^{l}\rrbracket\cup\llbracket\sigma 10^{l}\rrbracket$
  • $R_{e}$ is satisfied since $f(x)\uparrow$ and $f$ is undefined in $\llbracket\sigma 00^{l}\rrbracket\cup\llbracket\sigma 10^{l}\rrbracket$.
In the meantime the remaining branches above $\sigma$ believe Case 1 and proceed to other requirements. If $\sigma 00^{l}$, $\sigma 10^{l}$ have the wrong guess there will be a stage where this belief is seen to be wrong. At this point, although $\hat{f}$ will have been defined above the remaining branches out of $\sigma$, we will extend the latter appropriately and reconsider $R_{e}$ on these newly added leaves.
Markers. The assignment of requirement $R_{e}$ to a $\llbracket\sigma\rrbracket$ is done by placing an $e$-marker on node $\sigma$ as elaborated above. The marker acts if the $\rho$ in (7) is found. At each stage we keep a record of the status of each requirement on each leaf node: initially $\lambda$ is $e$-unattended for each $e$ and
  • when $\sigma$ receives an $e$-marker, $\sigma 00^{l}$, $\sigma 10^{l}$ become $e$-waiting and the other branches are temporarily $e$-satisfied
  • when this $e$-marker acts, the $e$-waiting branch $\sigma(1-i)0^{l}$ becomes $e$-satisfied and the other $e$-waiting branch along with the temporarily $e$-satisfied nodes extending $\sigma$ become $e$-unattended
  • in all other cases the status of the leaf node is the same as its parent.
A node is inactive if it is $e$-waiting for some $e$, indicating that no definition of $\hat{f}$ is permitted above them under this status. We ensure that
$f(x)\uparrow\iff\textrm{$x\upharpoonright_{n}$ stays inactive forever for some$n$.}$
Positive domain. As we elaborated, if an $e$-marker on $\sigma$ never acts:
  • $f$ becomes undefined on measure $\leq 2^{-l}$ relative to $\llbracket\sigma\rrbracket$
  • no other $e$-markers are placed above $\sigma$ until the $e$-marker on $\sigma$ acts
so the set of strings with $e$-marker that never acts is prefix-free. Thus each $R_{e}$ can only cause $f$ undefined on measure $\leq 2^{-l}$. By letting $l=e+k$ for $k:=2$:
$\mu(\mathsf{dom}(f))\geq 1-\sum_{e}2^{-e-k}=1-2^{1-k}>0.$
Injective. We make $f$ injective by ensuring that at any stage $s$:
  1. (i)
    for any active leaf nodes $\sigma\ |\ \tau$ we have $\hat{f}_{s}(\sigma)\ |\ \hat{f}_{s}(\tau)$,
  2. (ii)
    if there is an $e$-marker on $\sigma$ then $f_{s+1}(\llbracket\sigma\rrbracket)\subseteq f_{s}(\llbracket\sigma\rrbracket)$ where
    $f_{s}(\llbracket\sigma\rrbracket):=\bigcup\left\{\llbracket\hat{f}_{s}(\sigma\tau)\rrbracket:\sigma\tau\text{ is an active leaf node of }\hat{f}_{s}\right\}$
    so $f(\llbracket\sigma\rrbracket)\subseteq f_{s}(\llbracket\sigma\rrbracket)$.
To this end we often extend $\hat{f}_{s}$ to $\rho$ naturally, which means that we find the unique leaf node $\sigma\prec\rho$ in $\hat{f}_{s}$ and let $\hat{f}_{s+1}(\sigma\tau)=\hat{f}_{s}(\sigma)\tau$ for all $\tau$ such that $\sigma\tau\preceq\rho$. Note that (i), (ii) are preserved under natural extensions.
4.2 Construction and verification
We divide the stages for the construction to 3 types:
  • initialization stages $3\langle e,t\rangle$ where we put new $e$-markers on $e$-unattended nodes and do the initialization from them
  • activation stages $13\langle e,t\rangle+$ 1 where $e$-markers requiring action act
  • expansion stages $23t+$ 2 where we expand $f$ at its active leaf nodes.
Construction. Let $k=2$, $\hat{f}_{0}(\lambda)=\lambda$ and $\lambda$ be $e$-unattended for each $e$.
Stage $s=3\langle e,t\rangle$: for each active $e$-unattended leaf node $\sigma$:
  • put an $e$-marker on $\sigma$
  • for $\tau=\sigma 00^{e+k}$ and $\tau=\sigma 10^{e+k}$ set $\hat{f}_{s+1}(\sigma\tau)=\hat{f}_{s}(\sigma)$, define $\hat{f}_{s+1}$ on strings between $\sigma$ and $\sigma\tau$, and set $\sigma\tau$ $e$-waiting;
  • for all other $\tau$ with $|\tau|=e+k+1$ naturally extend $\hat{f}_{s}$ to $\sigma\tau$ and set $\sigma\tau$ temporarily $e$-satisfied.
Stage $s=3\langle e,t\rangle+1$: for each $e$-marker on $\sigma$ with $|\sigma|=n$ that has not acted, find $\tau$ with $\hat{f}_{s}(\sigma\tau)\downarrow=\hat{f}_{s}(\sigma)\rho$ and $\Phi_{e}(\rho)(n)[s]\downarrow=i$. If $\tau$ does not exist proceed to the next stage. Otherwise we may assume that $\sigma\tau$ is a leaf node of $\hat{f}_{s}$, since otherwise we can replace $\tau$ with an extension of it on leaf nodes. We then set:
  • $j=1-i$, $\hat{f}_{s+1}(\sigma j0^{e+k})=\hat{f}_{s}(\sigma)\rho 00$ and $\sigma j0^{e+k}$ be $e$-satisfied
  • $\sigma i0^{e+k}$ be $e$-unattended, $f_{s+1}(\sigma\tau)=f_{s}(\sigma)\rho 1$, $\hat{f}_{s+1}(\sigma i0^{e+k})=\hat{f}_{s}(\sigma)\rho 01$
  • all temporarily $e$-satisfied leaf nodes of $\hat{f}_{s}$ extending $\sigma$ be $e$-unattended
and declare that this $e$-marker has acted.222suffixing $\hat{f}_{s}(\sigma)\rho$ with $00,1,01$ serves satisfying (i), (ii) above.
Stage $s=3\langle e,t\rangle+2$: for each active leaf node $\sigma$ naturally extend $\hat{f}_{s}$ to $\sigma 0,\sigma 1$. This ensures $f(x)\downarrow$ when $x$ does not have prefixes that stay inactive forever.
Verification.
As we discussed in §4.1 the domain of $f$ is positive.
By induction on $s$ we get (i), (ii) from §4.1.
Lemma 4.3.
The function $f$ is injective.
Proof..
Assuming $x\neq y$, $f(x),f(y)\downarrow$ we show $f(x)\neq f(y)$.
Let $\chi_{s}$ be the unique leaf node of $\hat{f}_{s}$ with $\chi_{s}\prec x$ and similarly for $\gamma_{s}\prec y$.
Since $x\neq y$ for sufficiently large $s$ we have $\chi_{s}\ |\ \gamma_{s}$. So if $\chi_{s}$ and $\gamma_{s}$ are both active at $s$ by (i) we get $\hat{f}_{s}(\chi_{s})\ |\ \hat{f}_{s}(\gamma_{s})$ so $f(x)\neq f(y)$. For this reason we may assume that at most one of $\chi_{s}$, $\gamma_{s}$ is active at sufficiently large $s$.
Without loss of generality pick $s$ such that $\chi_{s}$ is active and $\gamma_{s}$ is inactive.
Since $f(y)\downarrow$ there exists $t>s$ such that
$\chi_{t}\succ\chi_{s}$ is inactive and $\gamma_{t}=\gamma_{s}$ is active.
Let $\sigma$ be the node with an $e$-marker that made $\chi_{t}$ inactive so
  • $\chi_{s}\preceq\sigma\prec\chi_{t}$ and $\chi_{s}\ |\ \gamma_{s}$, $\tau\ |\ \gamma_{t}$ for any leaf node $\tau\succ\sigma$
  • by (i) we have $\hat{f}_{t}(\tau)\ |\ \hat{f}_{t}(\gamma_{t})$ so $f_{t}(\llbracket\sigma\rrbracket)\cap\llbracket\hat{f}_{t}(\gamma_{t})\rrbracket=\emptyset$.
By (ii) we have $f(x)\in f(\llbracket\sigma\rrbracket)\subset f_{t}(\llbracket\sigma\rrbracket)$ and $f(y)\in\llbracket\hat{f}_{t}(\gamma_{t})\rrbracket$, so $f(x)\neq f(y)$. ∎
It remains to show that if $g=\Phi_{e}$ for some $e$ then $g$ inverts $f$ on measure 00.
For a contradiction assume otherwise so $\mu(L)>0$ where
$L=\{x:\Phi_{e}(f(x))=x\}.$
By Lebesgue’s density theorem, there is $x_{0}\in L$ with density $1$ in $L$. For each $s$ let $\chi_{s}$ be the unique leaf node of $\hat{f}_{s}$ with $\chi_{s}\prec x_{0}$. Then:
  • if $\chi_{s}$ is $e$-satisfied the $e$-status of $\chi_{t}$ remains $e$-satisfied forever;
  • if $\chi_{s}$ is $e$-waiting, since $f(x_{0})\downarrow$ there must be $t>s$ where $\chi_{t}=\chi_{s}$ becomes $e$-satisfied or $e$-unattended;
  • if $\chi_{s}$ is $e$-unattended there must be $t>s$ where an $e$-marker is put on some $\chi_{t}$ and $\chi_{t+1}$ is $e$-waiting or temporarily $e$-satisfied;
  • if $\chi_{s}$ is temporarily $e$-satisfied, the $e$-status of $\chi_{t}$ might remain temporarily $e$-satisfied forever, or might become $e$-unattended for some $t>s$.
Therefore there are three cases for the status of $\chi_{s}$ as $s\to\infty$.
Case 1. $\chi_{s}$ is $e$-satisfied for all sufficiently large $s$. Then $R_{e}$ is satisfied on all $x\in\chi_{s}$ so $R_{e}$ is satisfied for $x_{0}$, contradicting $x_{0}\in L$.
Case 2. $\chi_{s}$ is temporarily $e$-satisfied for all sufficiently large $s$. Then the $e$-marker on some $\sigma\prec\chi_{s}$ never acts so the corresponding $\tau,\rho$ are never found and $R_{e}$ is satisfied in $\llbracket\sigma\rrbracket$. Since $x_{0}\in\llbracket\sigma\rrbracket$ this contradicts $x_{0}\in L$.
Case 3: the $e$-status of $\chi_{s}$ as $s\to\infty$ loops indefinitely through
  1.     (a)
    $e$-unattended
  2.     (b)
    $e$-waiting or temporarily $e$-satisfied
  3.     (c)
    $e$-unattended
Transition (a)$\to$(b) is caused by an $e$-marker on some $\sigma$ which will act when transition (b)$\to$(c) occurs. Each such action produces a set of relative measure $2^{-e-k-1}$ where $R_{e}$ is satisfied, so the density of $x_{0}$ in $L$ is $\leq 1-2^{-e-k-1}$, contradicting the hypothesis that $x_{0}$ has density $1$ in $L$.
We conclude that $\mu(L)=0$ so all requirements are met.